OpenStudy (kainui):

I have a trick question for you, and a medal if you answer it right What two vectors have a dot product equal to e^(ix)?

OpenStudy (amistre64):

1,0 and a,k will dot to: a

OpenStudy (amistre64):

ix = 9 if my roman numeral classes are remembered correctly

OpenStudy (kainui):

Hahaha clever, but not what I was looking for. If no one gives a better answer you have it. Also, I should mention: Both vectors are not the same vector, they both have a length of 1, but neither of them have an entry that is 0 or e^(ix).

OpenStudy (amistre64):

\[|a||b|~cos\alpha=a\cdot b\] \[|a||b|~cos\alpha=e^{ix}\] if a and b are unit length for simplicity \[cos\alpha=e^{ix}\]

OpenStudy (amistre64):

e^(ix) seem to recall a trig definition for imaginary e stuff

terenzreignz (terenzreignz):

what about (cosx , sinx) and (1,i) ?

OpenStudy (kainui):

Yeah haha that's it, not really that tricky unfortunately.

terenzreignz (terenzreignz):

You threw me off >:( (1,i) doesn't have unit length

OpenStudy (kainui):

Yeah you're totally right, and I'm totally wrong haha.

terenzreignz (terenzreignz):

^^_ oh well

OpenStudy (kainui):

(1+i)(1-i)=1+1=2 != 1 what am I doing with my life, I don't even know.